# AOMGen: Photoreal, Physics-Consistent Demonstration Generation for Articulated Object Manipulation

### v1

@@ 1 Introduction
Robotic manipulation models based on Vision-Language-Action (VLA), which integrate visual perception, language understanding, and action execution, have demonstrated remarkable generalization capabilities across various tasks [intelligence2025pi05visionlanguageactionmodelopenworld, gr00tn1_2025, kim24openvla]. Achieving such capabilities typically requires substantial amounts of high-quality robot manipulation demonstration data collected from real-world interactions [open_x_embodiment_rt_x_2023, wu2025robomind, jones24fuse]. In particular, fine-grained manipulation tasks involving articulated objects rely more heavily on diverse and precise data compared to other tasks, as their complex kinematic constraints require richer demonstrations for learning effective control policies [kerr2024rsrd, xue2025demogen]. Nevertheless, collecting sufficient high-quality demonstration data remains expensive, labor-intensive, and inherently limited in scenario coverage, making it challenging to obtain adequate data for these tasks.
@@
To overcome the limitations posed by the inefficient collection of real-world robot manipulation data, existing research has primarily explored physics-based simulation and video-driven world modeling as two mainstream strategies. Physics-based simulation platforms [makoviychuk2021isaacgymhighperformance, Genesis, Wang_2025_CVPR] efficiently generate large, state–action aligned demonstrations, reducing manual annotation and hardware costs [li24simpler, long2025surveylearningembodiedintelligence], which have enabled several methods to achieve promising results by leveraging simulated data. However, simulation environments still exhibit significant gaps in visual realism compared to real-world scenes, posing substantial challenges for Sim-to-Real transfer [li2024robogsimreal2sim2realroboticgaussian, jia2025discoverseefficientrobotsimulation]. Recent video world model approaches enhance data realism by directly learning from large-scale real-world videos. [jang2025dreamgen, zhu2025uwm, zhou2024robodreamerlearningcompositionalworld, yang2025orv] Compared with purely simulation-based methods, they offer higher visual fidelity and diversity while reducing reliance on handcrafted simulation assets. However, these models often provide insufficient supervision over physical realism and action executability, leading to physically inconsistent interactions. Hence, for generating visually realistic and physically stable demonstration data, DemoGen [xue2025demogen] and R2RGen [xu2025r2rgenrealtoreal3ddata] synthesize additional robot manipulation demonstrations from limited real-world examples, including spatially augmented end-effector trajectories and 3D visual observations. However, they have several critical limitations: (1) limited to simple grasping and placement, unable to handle fine-grained articulated manipulation; (2) fixed object appearances and geometries, restricting generalization to novel objects or poses; and (3) single-view inputs, reducing visual realism in multi-view observations. These limitations underscore the need for methods capable of generating more visually realistic and physically consistent demonstrations for articulated object manipulation tasks.
@@ 2.1 Robot Manipulation Data Generation
MimicGen and its extensions [mandlekar2023mimicgendatagenerationscalable, hoque2024intervengeninterventionaldatageneration, jiang2025dexmimicgenautomateddatageneration] reconfigure the collected demonstration data to synthesize corresponding manipulation data. This provides a good generalization solution for seen objects; however, it fails to transfer to manipulation cognition for unseen objects.
Another line of work leverages LLM and VLM to generate manipulation data from a single image [chen2023genaug, jang2025dreamgen, zhu2025uwm, zhou2024robodreamerlearningcompositionalworld, yang2025orv] or generalize existing observation data [zhou2024autonomous]. These methods have been effective in visual authenticity and efficiency, but due to the uncontrollability of the generative models, the physical authenticity of the generated data cannot be guaranteed. Therefore, AOMGen reconstructs and generalizes one real-world collected data to ensure both its visual and physical performance are excellent.
@@ 2.2 Gaussian Splatting Reconstruction & Editing
3D Gaussian Splatting, as an explicit radiance field representation, possesses real-time rendering and interpretable editing capabilities. Some Real-to-Sim-Real pipelines leverage 3DGS to reconstruct scenes and reduce visual discrepancy with the real world [wu2025rlgsbridge3dgaussiansplatting, lou2024robogsphysicsconsistentspatialtemporal, han2025re3simgeneratinghighfidelitysimulation, torne2024reconcilingrealitysimulationrealtosimtoreal]. Furthermore, to achieve physically realistic interactions, some works directly edit Gaussian points to reconstruct manipulation trajectories consistent with real-world demonstrations from humans [kerr2024rsrd, yu2025real2render2realscalingrobotdata] or robots [qureshi2024splatsimzeroshotsim2realtransfer, robosplat]. Specially, Sizhe Yang et al. [robosplat] propose RobotSplat that leverages 3DGS to generate novel demonstrations. But RobotSplat fails to generalize manipulation data for articulated objects.
Unlike the robot arm trajectory, the motion trajectory of the target object has no real recorded data for direct reference. Justin Kerr et al. [kerr2024rsrd] provide a part-level motion recovery method based on DINO features, but how to accurately transfer it to an unseen object remains an open problem. Therefore, we propose a motion transfer method that leverages the real robot arm trajectories to supervise Gaussian editing, thereby generating physically realistic and generalized robotic manipulation data.
On the other hand, Gaussian Editing [chen2023gaussianeditor] and Inpainting [huang20253d] is the key factor that determines the visual quality of the generated data. How to correctly segment the Gaussian points of the object to be replaced and enhance the material appearance of the new object to make it more consistent with the real-world environment will determine the rendering realism after Gaussian editing.
@@ 2.3 Articulated Objects Model
Compared to rigid objects, the motion of articulated objects is constrained by the joints, making their motion patterns more complex than simple transformations in the SO(3) space. Recent works [10610171, liu2025videoartgsbuildingdigitaltwins] model articulated objects by using multi-view RGB images captured before and after interaction, thereby obtaining the geometric structure of articulated objects (such as the pose of the joints and different parts). However, the above methods fail in the absence of enough visual observations after interaction. Therefore, we propose a training-free articulated object modeling approach that explicitly captures a series of features by analyzing its geometric structure and incorporating robotic arm manipulation data.
@@ 4.1.1 Part-level Segmented 3DGS
Fine-grained segmentation is a necessary condition for editable GS scenes. Based on SAGA [cen2023saga], each Gaussian point is attached with a feature $f^{D}$ of dimensions $D$, which is learned from multi-view 2D masks extracted by SAM2 [ravi2024sam2]. By adjusting the segmentation granularity of SAM2, part-level image masks can be obtained.
@@ 4.1.2 Coordinate Alignment
The current GS coordinate system should be aligned with the real-world coordinate frame to ensure consistency in real-world, simulation and GS scenes. Given the robot URDF in the real-world coordinate frame, we can sample point clouds $P_{urdf}$ from the mesh surface. Using Iterative Closest Point (ICP) [zhang2021fast] on $P_{urdf}$ and $P_{robot}$, which are the robot points segmented from $G$, we can get the transformation matrix $T_{gs\rightarrow real}$ and the transformed 3DGS scene $G^{{}^{\prime}}$.
@@ 4.3 Articulated Object Replacement
Providing the same category of articulated object $AO_{new}$, we propose a novel process to process $AO_{new}$, including the geometry and appearance, as illustrated in the right part of Fig 2. After processing the given articulated object asset, we convert the USD file into a 3DGS representation [scolari2025mesh2splat]. Then, the motion simulation of the new movable part can be obtained using the method from Section 4.2.4. Moreover, to expand the generalization boundary, the pose of the replacement object can be arbitrarily modified.
@@ 4.3.1 Physical Interaction Adaptation
To get the accurate contact point of $AO_{new}$, a contact point mapping method inspired by NOCS [Wang_2019_CVPR] is proposed. Specifically, the point clouds $Part_{move}$ and $Part_{move}^{new}$ (obtained by sampling points from the mesh surface) are normalized to a unit cube space. Then, the relative position of $PC_{move}$ in the normalized space is projected onto the surface of the movable part of the newly normalized articulated object. Finally, by applying denormalization, the contact point information of the new articulated object $PC_{map}$ is obtained.
@@ 4.3.2 Visual Enhancement
At the same time, the visual performance of the replaced object is also crucial for the reality of the generated data. We apply the lighting changes from the real scene to the USD. First, we use DiffusionLight [Phongthawee2023DiffusionLight] to extract the lighting from the real environment. Then, bake the ambient lighting onto the object’s material in Blender. Moreover, Gaussian Inpainting is applied to handle the Gaussian holes caused by object replacement.
@@ 5.1 Experiments Setup
To demonstrate the robustness of AOMGen in handling a variety of articulated objects, we select three different simulation assets with rotational joint, including microwave ovens, tool box and computer, and two assets with prismatic joint, including drawer and cabinet from ArtVIP [jin2025artviparticulateddigitalassets] for substitution. During pose generalization, the translation range of the object is [0.05m,0.3m][0.05m,0.05m][-0.05m,0.3m]*[-0.05m,0.05m], and the rotation range is $45^{\circ},45^{\circ}]$.
The entire model training is carried out on an NVIDIA RTX4090 GPU. Specially, the model $\pi_{0.5}$ [intelligence2025pi05visionlanguageactionmodelopenworld] is fine-tuned using LoRA, with a batch size of 16 and a learning rate of $5\times 10^{-5}$, employing cosine decay with a 10K step warm-up. The training consists of 30,000 steps, and the optimizer used is AdamW with gradient clipping set to 1.0. The action prediction horizon is set to 5 steps, and precision is configured with bfloat16 (frozen) and float32 (trainable). The model OpenVLA [kim24openvla] is fine-tuned by LoRA using 50,000 steps, enabled with a rank of 32. Moreover, we use IsaacSim as our simulation platform.
@@ 5.3 VLA Training
For each replacement object, during the inference phase, we import the simulation assets of the robotic arm, the articulated object, and the table. To reduce the gap between the simulation environment and the real world, we add a ’background wall’ in the simulation environment and apply 3DGS render images, which have already undergone coordinate frame alignment, as textures on the background wall to simulate the real operating scene. To ensure the validity of the experimental results, we conduct experiments on two different models: $\pi_{0.5}$ [intelligence2025pi05visionlanguageactionmodelopenworld] and OpenVLA [kim24openvla]. What’s more, each evaluation is conducted with 30 trials.

### v2

@@ 1 Introduction
Robotic manipulation models based on Vision-Language-Action (VLA), which integrate visual perception, language understanding, and action execution, have demonstrated remarkable generalization capabilities across various tasks [9, 24, 16]. Achieving such capabilities typically requires substantial amounts of high-quality robot manipulation demonstration data collected from real-world interactions [5, 33, 14]. In particular, fine-grained manipulation tasks involving articulated objects rely more heavily on diverse and precise data compared to other tasks, as their complex kinematic constraints require richer demonstrations for learning effective control policies [15, 36]. Nevertheless, collecting sufficient high-quality demonstration data remains expensive, labor-intensive, and inherently limited in scenario coverage, making it challenging to obtain adequate data for these tasks.
@@
To overcome the limitations posed by the inefficient collection of real-world robot manipulation data, existing research has primarily explored physics-based simulation and video-driven world modeling as two mainstream strategies. Physics-based simulation platforms [22, 1, 32] efficiently generate large, state–action aligned demonstrations, reducing manual annotation and hardware costs [18, 20], which have enabled several methods to achieve promising results by leveraging simulated data. However, simulation environments still exhibit significant gaps in visual realism compared to real-world scenes, posing substantial challenges for Sim-to-Real transfer [17, 11]. Recent video world model approaches enhance data realism by directly learning from large-scale real-world videos. [10, 43, 41, 38] Compared with purely simulation-based methods, they offer higher visual fidelity and diversity while reducing reliance on handcrafted simulation assets. However, these models often provide insufficient supervision over physical realism and action executability, leading to physically inconsistent interactions. Hence, for generating visually realistic and physically stable demonstration data, DemoGen [36] and R2RGen [35] synthesize additional robot manipulation demonstrations from limited real-world examples, including spatially augmented end-effector trajectories and 3D visual observations. However, they have several critical limitations: (1) limited to simple grasping and placement, unable to handle fine-grained articulated manipulation; (2) fixed object appearances and geometries, restricting generalization to novel objects or poses; and (3) single-view inputs, reducing visual realism in multi-view observations. These limitations underscore the need for methods capable of generating more visually realistic and physically consistent demonstrations for articulated object manipulation tasks.
@@ 2.1 Robot Manipulation Data Generation
MimicGen and its extensions [23, 7, 12] reconfigure the collected demonstration data to synthesize corresponding manipulation data. This provides a good generalization solution for seen objects; however, it fails to transfer to manipulation cognition for unseen objects.
Another line of work leverages LLM and VLM to generate manipulation data from a single image [4, 10, 43, 41, 38] or generalize existing observation data [42]. These methods have been effective in visual authenticity and efficiency, but due to the uncontrollability of the generative models, the physical authenticity of the generated data cannot be guaranteed. Therefore, AOMGen reconstructs and generalizes one real-world collected data to ensure both its visual and physical performance are excellent.
@@ 2.2 Gaussian Splatting Reconstruction & Editing
3D Gaussian Splatting, as an explicit radiance field representation, possesses real-time rendering and interpretable editing capabilities. Some Real-to-Sim-Real pipelines leverage 3DGS to reconstruct scenes and reduce visual discrepancy with the real world [34, 21, 6, 29]. Furthermore, to achieve physically realistic interactions, some works directly edit Gaussian points to reconstruct manipulation trajectories consistent with real-world demonstrations from humans [15, 39] or robots [26, 37]. Specially, Sizhe Yang et al. [37] propose RobotSplat that leverages 3DGS to generate novel demonstrations. But RobotSplat fails to generalize manipulation data for articulated objects.
Unlike the robot arm trajectory, the motion trajectory of the target object has no real recorded data for direct reference. Justin Kerr et al. [15] provide a part-level motion recovery method based on DINO features, but how to accurately transfer it to an unseen object remains an open problem. Therefore, we propose a motion transfer method that leverages the real robot arm trajectories to supervise Gaussian editing, thereby generating physically realistic and generalized robotic manipulation data.
On the other hand, Gaussian Editing [3] and Inpainting [8] is the key factor that determines the visual quality of the generated data. How to correctly segment the Gaussian points of the object to be replaced and enhance the material appearance of the new object to make it more consistent with the real-world environment will determine the rendering realism after Gaussian editing.
@@ 2.3 Articulated Objects Model
Compared to rigid objects, the motion of articulated objects is constrained by the joints, making their motion patterns more complex than simple transformations in the SO(3) space. Recent works [30, 19] model articulated objects by using multi-view RGB images captured before and after interaction, thereby obtaining the geometric structure of articulated objects (such as the pose of the joints and different parts). However, the above methods fail in the absence of enough visual observations after interaction. Therefore, we propose a training-free articulated object modeling approach that explicitly captures a series of features by analyzing its geometric structure and incorporating robotic arm manipulation data.
@@ 4.1.1 Part-level Segmented 3DGS
Fine-grained segmentation is a necessary condition for editable GS scenes. Based on SAGA [2], each Gaussian point is attached with a feature $f^{D}$ of dimensions $D$, which is learned from multi-view 2D masks extracted by SAM2 [27]. By adjusting the segmentation granularity of SAM2, part-level image masks can be obtained.
@@ 4.1.2 Coordinate Alignment
The current GS coordinate system should be aligned with the real-world coordinate frame to ensure consistency in real-world, simulation and GS scenes. Given the robot URDF in the real-world coordinate frame, we can sample point clouds $P_{urdf}$ from the mesh surface. Using Iterative Closest Point (ICP) [40] on $P_{urdf}$ and $P_{robot}$, which are the robot points segmented from $G$, we can get the transformation matrix $T_{gs\rightarrow real}$ and the transformed 3DGS scene $G^{{}^{\prime}}$.
@@ 4.3 Articulated Object Replacement
Providing the same category of articulated object $AO_{new}$, we propose a novel process to process $AO_{new}$, including the geometry and appearance, as illustrated in the right part of Fig 2. After processing the given articulated object asset, we convert the USD file into a 3DGS representation [28]. Then, the motion simulation of the new movable part can be obtained using the method from Section 4.2.4. Moreover, to expand the generalization boundary, the pose of the replacement object can be arbitrarily modified.
@@ 4.3.1 Physical Interaction Adaptation
To get the accurate contact point of $AO_{new}$, a contact point mapping method inspired by NOCS [31] is proposed. Specifically, the point clouds $Part_{move}$ and $Part_{move}^{new}$ (obtained by sampling points from the mesh surface) are normalized to a unit cube space. Then, the relative position of $PC_{move}$ in the normalized space is projected onto the surface of the movable part of the newly normalized articulated object. Finally, by applying denormalization, the contact point information of the new articulated object $PC_{map}$ is obtained.
@@ 4.3.2 Visual Enhancement
At the same time, the visual performance of the replaced object is also crucial for the reality of the generated data. We apply the lighting changes from the real scene to the USD. First, we use DiffusionLight [25] to extract the lighting from the real environment. Then, bake the ambient lighting onto the object’s material in Blender. Moreover, Gaussian Inpainting is applied to handle the Gaussian holes caused by object replacement.
@@ 5.1 Experiments Setup
To demonstrate the robustness of AOMGen in handling a variety of articulated objects, we select three different simulation assets with rotational joint, including microwave ovens, tool box and computer, and two assets with prismatic joint, including drawer and cabinet from ArtVIP [13] for substitution. During pose generalization, the translation range of the object is [0.05m,0.3m][0.05m,0.05m][-0.05m,0.3m]*[-0.05m,0.05m], and the rotation range is $45^{\circ},45^{\circ}]$.
The entire model training is carried out on an NVIDIA RTX4090 GPU. Specially, the model $\pi_{0.5}$ [9] is fine-tuned using LoRA, with a batch size of 16 and a learning rate of $5\times 10^{-5}$, employing cosine decay with a 10K step warm-up. The training consists of 30,000 steps, and the optimizer used is AdamW with gradient clipping set to 1.0. The action prediction horizon is set to 5 steps, and precision is configured with bfloat16 (frozen) and float32 (trainable). The model OpenVLA [16] is fine-tuned by LoRA using 50,000 steps, enabled with a rank of 32. Moreover, we use IsaacSim as our simulation platform.
@@ 5.3 VLA Training
For each replacement object, during the inference phase, we import the simulation assets of the robotic arm, the articulated object, and the table. To reduce the gap between the simulation environment and the real world, we add a ’background wall’ in the simulation environment and apply 3DGS render images, which have already undergone coordinate frame alignment, as textures on the background wall to simulate the real operating scene. To ensure the validity of the experimental results, we conduct experiments on two different models: $\pi_{0.5}$ [9] and OpenVLA [16]. What’s more, each evaluation is conducted with 30 trials.